# Trace field degrees in the Torelli group

### v1

@@ 5. Proof of 3
The goal of this section is to realise every positive integer $d\leq 3g-3$ as the multicurve intersection degree of a pair of multicurves $\alpha,\beta\subset S$ on $S_{g}$ for $g\geq 3$ in such a way that the multicurves $\alpha$ and $\beta$ consist of components that are separating or that come in bounding pairs, where for each bounding pair one of the curves is a component of $\alpha$ and the other is a component of $\beta$. 4 then provides 3 in the case $g\geq 3$. For the case $g=2$ we note that the statement is proved for $d=3g-3=3$ in Section 4, and for $d\leq 2=g$ it is proved in [LL24].
We start with the maximal degree $33g-$ 3 and then discuss how to adapt the construction in order to realise smaller degrees.
5.1. Multicurve intersection degree $33g-$ 3
We start by realising, on the surface of genus $g\geq 2$ with one boundary component, a pair of filling multicurves $\alpha$ and $\beta$ such that $\chi_{XX^{\top}}\in\mathbb{Z}[t]$ is irreducible and of degree $23g-$ 2, in such a way that the multicurves $\alpha$ and $\beta$ consist of components that are separating or that come in bounding pairs, where for each bounding pair one of the curves is a component of $\alpha$ and the other is a component of $\beta$. The construction is done by induction on the genus $g\geq 2$.
@@ 5.1. Multicurve intersection degree $33g-$ 3
We consider the two multicurves $\alpha$ and $\beta$ shown in Figure 3.
@@
Figure 3. Two multicurves $\alpha$ and $\beta$ on the surface of genus two with one boundary component. One component of $\beta$ has $y$ parallel copies.
We first note that the components $\alpha_{1}$ and $\alpha_{3}$ are separating. Furthermore, the components $\alpha_{2}$ and $\alpha_{4}$ have their counterparts in the multicurve $\beta$ with which they each form a bounding pair. Finally, the component of $\beta$ of which there are $y$ parallel copies and the component of $\beta$ drawn in light blue in Figure 3 are separating.
@@ 5.1.2. For $g>2$ and one boundary component
For the inductive step, let $g\geq 2$. Assume we have constructed on the surface of genus $g$ with one boundary component a pair of multicurves $\alpha^{\prime},\beta^{\prime}$ such that the characteristic polynomial $\chi_{XX^{\top}}\in\mathbb{Z}[t]$ is irreducible and of degree $23g-$ 2, in such a way that the multicurves $\alpha$ and $\beta$ consist of components that are separating or that come in bounding pairs, where for each bounding pair one of the curves is a component of $\alpha$ and the other is a component of $\beta$. Further, assume that $\alpha_{1}^{\prime}$ is a simple closed curve that encircles all the handles of the surface, except for the rightmost one. Then, we take such a model surface and glue to its boundary a surface of genus one with two boundary components, as shown in Figure 4, and add two new curves $\alpha_{0}$ and $\beta_{0}$ to the multicurves. The curve $\alpha_{0}$ encircles all the handles of the newly formed surface, except for the rightmost one, and the curve $\beta_{0}$ runs along the glued boundary components, and twice intersects $\alpha_{0}$ but no other component of $\alpha$, see Figure 5.
@@
$33g^{\prime}-2+3g^{\prime\prime}-2+1=3g-$ 3 .
@@
Figure 6. Two multicurves $\alpha$ and $\beta$ on the surface of genus three. There are to new components of $\alpha$ when compared to Figure 3: a nonseparating component (red) that we call $\alpha_{5}$ and a separating component (orange) that we call $\alpha_{6}$. Similarly, there are two new components of $\beta$: a separating component (blue) that we call $\beta_{5}$ and a nonseparating component (light blue) that we call $\beta_{6}$.
First add the curves $\alpha_{5}$ and $\beta_{5}$ with $y^{2}-29$ parallel copies. The resulting characteristic polynomials is irreducible for infinitely many choices of $y$ by 9. Repeat the same process with $\alpha_{6}$ and $\beta_{6}$ (adjusting the number of parallel copies of $\beta_{6}$ suitably), and we are done.
5.2. Multicurve intersection degrees $d<3g-3$
We now show how to modify our construction from Section 5.1 in order to realise multicurve intersection degrees smaller than the maximal multicurve intersection degree $33g-$ 3. We need new building blocks to construct our surfaces.
@@ 5.2. Multicurve intersection degrees $d<3g-3$
Figure 7. Two separating and filling curves $\alpha$ and $\beta$ on the surface of genus two with one boundary component.
Block 3. Take a surface as depicted in Figure 8. We denote the red multicurve by $\alpha$ and the blue multicurve by $\beta$. The multicurve $\alpha$ has $k+1$ separating components: one for each of the handles that separates the handle, and one that separates all the handles. We denote the component of $\alpha$ that separates all the handles of the surface in Figure 8 by $\alpha_{1}$, and we denote the other separating components of $\alpha$ by $\alpha_{2},\alpha_{4},\dots,\alpha_{2g-2}$ from left to right. Finally, the remaining nonseparating components of $\alpha$ are $\alpha_{3},\alpha_{5},\dots,\alpha_{2g-1}$ from left to right.
@@
Figure 8. A surface of genus $k$ with two boundary components, as well as two multicurves $\alpha$ (in red) and $\beta$ (in blue). The separating components of $\beta$ can have several parallel copies: the ones separating the handles have $y_{1},\dots,y_{k}$ copies, and the separating component in the middle has $y^{2}-4k-y_{1}-\cdots-y_{k}$ copies.
In this situation, we have
@@
[…]
5.2.1. Realising multicurve intersection degrees $33g-6\leq d<3g-$ 3.
Using a block of type 1 or 2 instead of our standard starting surface depicted in Figure 3, we can reduce the multicurve intersection degree by 1 or 3, respectively. Since we use such block on both sides of the surface in our construction, this gives the possibility to reduce the degree by any among the numbers 1,2,3,4 or 6. In particular, we can clearly realise the multicurve intersection degrees $43g-3,3g-$ 4 and $53g-$ 5. This argument works for $g\geq 4$.
In case of $g=3$, we need a separate argument. The idea is to copy our example of maximal degree from Figure 6, but leave out first $\alpha_{3}$ and then also $\alpha_{1}$. We start from the multicurves depicted in Figure 3 and drop $\alpha_{3}$. Letting $y=2$, we then get
$XX^{\top}=\begin{pmatrix}116&56&16\\ 56&28&8\\ 16&8&4\end{pmatrix},$
which has irreducible characteristic polynomial. We can now close off the surface by glueing a torus with one boundary component and add more components to $\alpha$ and $\beta$, in the same way as in Figure 6. The exact same argument we used to realise degree $6$ now yields degree $5$ instead.
In order to realise degree 4 for $g=3$, we note that if we start from the multicurves depicted in Figure 6 and drop the components $\alpha_{1},\alpha_{3},\alpha_{6}$ as well as $\beta_{5},\beta_{6}$, then the matrix $XX^{\top}$ for $\alpha_{5},\alpha_{2},\alpha_{4}$ is exactly the matrix as above:
$XX^{\top}=\begin{pmatrix}116&56&16\\ 56&28&8\\ 16&8&4\end{pmatrix},$
with irreducible characteristic polynomial. If we add back $\alpha_{6}$ and $\beta_{6}$ with $y^{2}-116$ parallel copies, the resulting characteristic polynomials is irreducible for infinitely many choices of $y$ by 9, realising degree 4. Note that all in all, we have dropped the components $\alpha_{1},\alpha_{3}$ and $\beta_{6}$, which are all separating. Therefore, we have not changed the fact that the multicurves $\alpha$ and $\beta$ consist of components that are separating or that come in bounding pairs, where for each bounding pair one of the curves is a component of $\alpha$ and the other is a component of $\beta$.
@@
We start by constructing a surface of genus $g-2$ with two boundary components, which we then close off in a second step.
Using surfaces of the type depicted in Figure 4 and applying the inductive step procedure, we can construct a surface of genus $g-2\geq 1$ and two boundary components, as well as filling multicurves $\alpha$ and $\beta$ with intersection degree $73(g-2)-1=3g-$ 7. Using at some point in the inductive procedure a block of type 3 of genus $k\leq g-2$, as depicted in Figure 8, we can reduce the degree by up to $62k-2\leq 2g-$ 6, realising multicurve intersection degrees from $g-1$ to $73g-$ 7 on the surface of genus $g-2$ with two boundary components. Now we close the surface, as depicted in Figure 9, adding the new components $\alpha_{0}$ and $\beta_{0}$ to the multicurves $\alpha$ and $\beta$, respectively.
@@
where $A$ is the matrix $XX^{\top}$ before adding the curves $\alpha_{0}$ and $\beta_{0}$. Since $a_{1}=4a$, we can set $\rho=y^{2}-a$ to have the top-left coefficient $64y^{2}$, which is exactly the form of the matrix in 9. Finishing the argument as usual, we can realise the multicurve intersection degrees $g\leq d\leq 3g-6$ for $g\geq 3$.
@@
For $2\leq d<g$, let us define $f=g-1-d$. We start with a surface block of type 3 of genus $g-2$, where we deleted all the components of $\alpha$ that are separating. We also remove the component of $\beta$ in the middle of Figure 8. Furthermore, we let the $f+1\leq g-2$ first of the parameters $y_{i}$ be equal to $1$. Then we close off the surface as in the previous case, adding one component $\alpha_{0}$ to $\alpha$ and one component $\beta_{0}$ to $\beta$, compare with Figure 9. Assume there are $\rho$ parallel copies of $\beta_{0}$. We get
@@
If all the $z_{i}$ are pairwise distinct, this polynomial is irreducible as a polynomial in $t,y$ by Lemma 10 in [LL24]. However, we chose that the first $f+1$ coefficients $y_{1},\dots,y_{f+1}$ are equal to $1$ and the other $y_{i}\neq 1$ and pairwise distinct. In particular, the polynomial $p(t,y)$ factors as $(t-4)^{f}\tilde{p}(t,y)$, where $\tilde{p}(t,y)$ is of degree $g-1-f=d$ in the variable $t$ and with pairwise distinct $z_{i}$. In particular, Lemma 10 in [LL24] implies that $\tilde{p}(t,y)\in\mathbb{Z}[t,y]$ is irreducible. Hilbert’s irreducibility theorem guarantees the existence of infinitely many positive specifications of $y$ such that the resulting polynomial is irreducible in $\mathbb{Z}[t]$.
@@
Finally, we end this section we a proof of 5.
@@
For every $g\geq 3$ and every integer $31\leq d\leq 3g-$ 3, we have constructed a pair of filling multicurves $\alpha$ and $\beta$, with a parameter $y$, such that $\chi_{XX^{\top}}(t,y)\in\mathbb{Z}[t,y]$ is irreducible. By Remark 7, we may run the same argument to show that also the polynomial $\chi_{XX^{\top}}(t^{2},y)\in\mathbb{Z}[t,y]$ is irreducible. By Hilbert’s irreducibility theorem, we find infinitely many specifications of $y$ such that $\chi_{XX^{\top}}(t^{2})\in\mathbb{Z}[t]$ is irreducible of degree $2d$. The leading eigenvalue $\mu$ of $XX^{\top}$ is a root of a characteristic polynomial $\chi_{XX^{\top}}(t)$, so $\chi_{XX^{\top}}(t^{2})\in\mathbb{Z}[t]$ is the minimal polynomial of $\sqrt{\mu}$. Hence, the multicurve bipartite degree of $\alpha$ and $\beta$ equals $2d$.

### v2

@@ 5. Proof of 3
The goal of this section is to realise every positive integer $d\leq 3g-3$ as the multicurve intersection degree of a pair of multicurves $\alpha,\beta\subset S$ on $S_{g}$ for $g\geq 3$ in such a way that the multicurves $\alpha$ and $\beta$ consist of components that are separating or that come in bounding pairs, where for each bounding pair one of the curves is a component of $\alpha$ and the other is a component of $\beta$. This guarantees that $T_{\alpha}\circ T_{\beta}^{-1}$ is an element of the Torelli group $\mathcal{I}(S_{g})$. Indeed, on the homology level, the action of a Dehn twist along a curve $\gamma$ is given by $\delta]\mapsto[\delta]+i(\delta,\gamma)[\gamma]$, where $i(\cdot,\cdot)$ is the algebraic intersection form on ${H}_{1}(S_{g};\mathbb{Z})$. It follows that a Dehn twist along a separating curve is en element of $\mathcal{I}(S_{g})$, but also each composition of Dehn twists $T_{\alpha_{i}}\circ T_{\beta_{j}}^{-1}$ for a bounding pair $\alpha_{i}$ and $\beta_{j}$. Since all the individual Dehn twists $T_{\alpha_{i}}$ commute, as well as all the individual Dehn twists $T_{\beta_{j}}$, it follows that if the multicurves $\alpha$ and $\beta$ consist of components that are separating or that come in bounding pairs, where for each bounding pair one of the curves is a component of $\alpha$ and the other is a component of $\beta$, the mapping class $T_{\alpha}\circ T_{\beta}^{-1}$ is an element of $\mathcal{I}(S_{g})$, and hence so is the mapping class $T_{\alpha}^{n}\circ T_{\beta}^{-n}$. 4 then provides 3 in the case $g\geq 3$. For the case $g=2$ we note that the statement is proved for d=3g3=3d=3g-3=3 in Section 4, and for $d\leq 2=g$ it is proved in [LL24].
We start with the maximal degree 3g33g-3 and then discuss how to adapt the construction in order to realise smaller degrees.
5.1. Multicurve intersection degree 3g33g-3
We start by realising, on the surface of genus $g\geq 2$ with one boundary component, a filling pair of multicurves $\alpha$ and $\beta$ such that $\chi_{XX^{\top}}\in\mathbb{Z}[t]$ is irreducible and of degree 3g23g-2, in such a way that the multicurves $\alpha$ and $\beta$ consist of components that are separating or that come in bounding pairs, where for each bounding pair one of the curves is a component of $\alpha$ and the other is a component of $\beta$. The construction is done by induction on the genus $g\geq 2$.
@@ 5.1. Multicurve intersection degree 3g33g-3
We consider the two multicurves $\alpha$ and $\beta$ shown in Figure 4, with $y$ parallel copies of the component $\beta_{1}$.
@@
of $\beta_{1}$
$\beta_{2}$
$\beta_{4}$
$\beta_{3}$
Figure 4. Two multicurves $\alpha$ and $\beta$ on the surface of genus two with one boundary component. There are $y$ parallel copies of the component $\beta_{1}$, and the separating component $\beta_{3}$ is drawn in light blue.
We first note that the components $\alpha_{1},\beta_{1},\alpha_{3}$ and $\beta_{3}$ are separating. Furthermore, the components $\alpha_{2}$ and $\alpha_{4}$ have their counterparts $\beta_{2}$ and $\beta_{4}$, respectively, in the multicurve $\beta$ with which they each form a bounding pair.
@@ 5.1.2. For $g>2$ and one boundary component
For the inductive step, let $g\geq 2$. Assume we have constructed on the surface of genus $g$ with one boundary component a pair of multicurves $\alpha^{\prime},\beta^{\prime}$ such that the characteristic polynomial $\chi_{XX^{\top}}\in\mathbb{Z}[t]$ is irreducible and of degree 3g23g-2, in such a way that the multicurves $\alpha$ and $\beta$ consist of components that are separating or that come in bounding pairs, where for each bounding pair one of the curves is a component of $\alpha$ and the other is a component of $\beta$. Further, assume that $\alpha_{1}^{\prime}$ is a simple closed curve that encircles all the handles of the surface, except for the rightmost one. Then, we take such a model surface and glue to its boundary a surface of genus one with two boundary components, as shown in Figure 5, and add two new curves $\alpha_{0}$ and $\beta_{0}$ to the multicurves. The curve $\alpha_{0}$ encircles all the handles of the newly formed surface, except for the rightmost one, and the curve $\beta_{0}$ runs along the glued boundary components, and twice intersects $\alpha_{0}$ but no other component of $\alpha$, see Figure 6.
@@
$\alpha_{1}^{\prime}$
$\alpha_{0}$
$\beta_{0}$
$\alpha^{\prime\prime}_{1}$
@@
$3g^{\prime}-2+3g^{\prime\prime}-2+1=3g-3.$
@@
2 copies
Figure 7. Two multicurves $\alpha$ and $\beta$ on the surface of genus three. There are to new components of $\alpha$ when compared to Figure 4: a nonseparating component (red) that we call $\alpha_{5}$ and a separating component (orange) that we call $\alpha_{6}$. Similarly, there are two new components of $\beta$: a separating component (blue) that we call $\beta_{5}$ and a nonseparating component (violet) that we call $\beta_{6}$.
We add the curves $\alpha_{5},\alpha_{6}$, two parallel copies of $\beta_{5}$ and finally the curve $\beta_{6}$. We note that $\beta_{5}$ and $\alpha_{6}$ are separating, and $\beta_{6}$ and $\alpha_{5}$ form a bounding pair. To read off the matrix $XX^{\top}$ for this pair of multicurves, we note two things. Firstly, the component $\alpha_{5}$ intersects the curves $\beta_{j}$ the same number of times as $\alpha_{1}$, except for the component $\beta_{5}$ of which there are two parallel copies. Secondly, the component $\alpha_{6}$ intersects each curve $\beta_{j}$ twice the number of times as $\alpha_{5}$, except for $\beta_{6}$. Using this, it is straightforward to calculate (ordering the curves as $\alpha_{6},\alpha_{5},\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}$)
$XX^{\top}=\begin{pmatrix}500&248&232&112&80&32\\ 248&124&116&56&40&16\\ 232&116&116&56&40&16\\ 112&56&56&28&20&8\\ 80&40&40&20&20&8\\ 32&16&16&8&8&4\end{pmatrix}$,
which is then checked to have irreducible characteristic polynomial.
5.2. Multicurve intersection degrees d<3g3d<3g-3
We now show how to modify our construction from Section 5.1 in order to realise multicurve intersection degrees smaller than the maximal multicurve intersection degree 3g33g-3. We need new building blocks to construct our surfaces.
@@ 5.2. Multicurve intersection degrees d<3g3d<3g-3
Figure 8. Two separating and filling curves $\alpha$ (red) and $\beta$ (blue) on the surface of genus two with one boundary component.
Block 3. Take a surface as depicted in Figure 9. We denote the red multicurve by $\alpha$ and the blue multicurve by $\beta$. The multicurve $\alpha$ has $k+1$ separating components: one for each of the handles that separates the handle, and one that separates all the handles. We denote the component of $\alpha$ that separates all the handles of the surface in Figure 9 by $\alpha_{1}$, and we denote the other separating components of $\alpha$ by $\alpha_{2},\alpha_{4},\dots,\alpha_{2k}$ from left to right. Finally, the remaining nonseparating components of $\alpha$ are $\alpha_{3},\alpha_{5},\dots,\alpha_{2k+1}$ from left to right.
@@
$\alpha_{1}$
$\alpha_{2}$
$\alpha_{2k}$
$\alpha_{3}$
$\alpha_{2k+1}$
$\beta_{1}$
$\beta_{k}$
$\beta_{0}$
Figure 9. A surface of genus $k$ with two boundary components, as well as two multicurves $\alpha$ (in red) and $\beta$ (in blue). The separating components of $\beta$ can have several parallel copies: the separating components $\beta_{1},\dots,\beta_{k}$ running through the handles have $y_{1},\dots,y_{k}$ copies, and the separating component $\beta_{0}$ to the right of the left-most handle has $y^{2}-k-4(y_{1}+\cdots+y_{k})$ copies.
The multicurve $\beta$ consists of a separating and a nonseparating component in each handle, as well as one separating curve $\beta_{0}$ to the right of the left-most handle. All the separating components of $\beta$ can have multiple parallel copies, see Figure 9. In this situation, we have
@@
If all the $y_{i}$ are chosen pairwise distinct,  Remark 12 guarantees that the polynomial $\chi_{XX^{\top}}(t,y)\in\mathbb{Z}[t][y]$ is irreducible. By Hilbert’s irreducibility theorem, there are infinitely many specifications of $y$ such that $y^{2}-k-4(y_{1}+\cdots+y_{k})>0$ and such that $\chi_{XX^{\top}}(t)\in\mathbb{Z}[t]$ is irreducible and of degree $2k+1$. Note that we can drop the separating components $\alpha_{2},\alpha_{4},\dots,\alpha_{2k}$ of $\alpha$ winding around one handle one by one in order to decrease the degree, reducing a 2-by-2 block to a 1-by-1 block, consisting of the coefficient $4y_{i}$, for each component dropped in this way. Again, if all the $y_{i}$ are chosen pairwise distinct, Remark 12 guarantees that the polynomial $\chi_{XX^{\top}}(t,y)\in\mathbb{Z}[t][y]$ is irreducible. We can in this way construct all degrees $k+1\leq d\leq 2k+1$ for the surface of genus $k$ and 2 boundary components.
5.2.1. Realising multicurve intersection degrees $3g-6\leq d<3g-3$.
Using a block of type 1 or 2 instead of our standard starting surface depicted in Figure 4, we can reduce the multicurve intersection degree by 1 or 3, respectively. Since we use such a block on both sides of the surface in our construction, this gives the possibility to reduce the degree by any among the numbers 1,2,3,4 or 6. In particular, we can clearly realise the multicurve intersection degrees 3g4,3g53g-4,3g-5 and 3g63g-6. This argument works for $g\geq 4$.
In case of $g=3$, we need a separate argument. The idea is to copy our example of maximal degree from Figure 7, but leave out first $\alpha_{1}$ and then also $\alpha_{3}$. If we drop $\alpha_{1}$, which amounts do deleting the third row and column, we obtain a new matrix $XX^{\top}$ which is directly verified to have an irreducible characteristic polynomial. If we now drop also $\alpha_{3}$, which amounts to deleting the second-last row and column, we obtain a new matrix $XX^{\top}$ which again is verified to have irreducible characteristic polynomial. Since we have only dropped separating components, we have not changed the fact that the multicurves $\alpha$ and $\beta$ consist of components that are separating or that come in bounding pairs, where for each bounding pair one of the curves is a component of $\alpha$ and the other is a component of $\beta$.
We have therefore realised $d=5,4$ for $g=3$. For $g=3$, the case $d=3$ equals the case d=3g6d=3g-6, which is treated below.
@@
We start by constructing a surface of genus g2g-2 with two boundary components, which we then close off in a second step.
Using surfaces of the type depicted in Figure 5 and applying the inductive step procedure illustrated in Figure 2, we can construct a surface of genus $g-2\geq 1$ and two boundary components, as well as filling multicurves $\alpha$ and $\beta$ with intersection degree 3(g2)1=3g73(g-2)-1=3g-7. The count of this intersection degree is exactly as in the non-Torelli case of Section 4, where we obtained degree 3g13g-1 on the closed surface of genus $g$ with 2 boundary components. Using at some point in the inductive procedure a block of type 3 of genus $k\leq g-2$, as depicted in Figure 9, we can reduce the degree by up to $2k-2\leq 2g-6$, realising multicurve intersection degrees from g1g-1 to 3g73g-7 on the surface of genus g2g-2 with two boundary components. Now we close the surface, as depicted in Figure 10, adding the new components $\alpha_{0}$ and $\beta_{0}$ to the multicurves $\alpha$ and $\beta$, respectively.
@@
where $A$ is the matrix $XX^{\top}$ before adding the curves $\alpha_{0}$ and $\beta_{0}$. Since $a_{1}=4a$, we can set $\rho=y^{2}-a$ to have the top-left coefficient $64y^{2}$, which is exactly the form of the matrix in 10. Finishing the argument as usual, we can realise the multicurve intersection degrees $g\leq d\leq 3g-6$ for $g\geq 3$.
@@
For $2\leq d<g$, let us define f=g1df=g-1-d. We start with a surface block of type 3 of genus g2g-2, where we deleted all the components $\alpha_{2},\alpha_{4},\dots,\alpha_{2k}$ of $\alpha$. We also remove the component $\beta_{0}$ of $\beta$, see Figure 9. Furthermore, we let the $f+1\leq g-2$ first of the parameters $y_{i}$ be equal to $1$. Then we close off the surface as in the previous case, adding one component $\alpha_{0}$ to $\alpha$ and one component $\beta_{0}$ to $\beta$, compare with Figure 10. Assume there are $\rho$ parallel copies of $\beta_{0}$. We get
@@
If all the $z_{i}$ are pairwise distinct, this polynomial is irreducible as a polynomial in $t,y$ by Lemma 10 in [LL24]. However, we chose that the first $f+1$ coefficients $y_{1},\dots,y_{f+1}$ are equal to 11 and the other $y_{i}\neq 1$ and pairwise distinct. In particular, the polynomial $p(t,y)$ factors as $(t-4)^{f}\tilde{p}(t,y)$, where $\tilde{p}(t,y)$ is of degree g1f=dg-1-f=d in the variable $t$ and with pairwise distinct $z_{i}$. In particular, Lemma 10 in [LL24] implies that $\tilde{p}(t,y)\in\mathbb{Z}[t,y]$ is irreducible. Hilbert’s irreducibility theorem guarantees the existence of infinitely many positive specifications of $y$ such that the resulting polynomial is irreducible in $\mathbb{Z}[t]$.
@@
Finally, we end this section with a proof of 5.
@@
For every $g\geq 3$ and every integer $1\leq d\leq 3g-3$, we have constructed a filling pair of multicurves $\alpha$ and $\beta$, with a parameter $y$, such that $\chi_{XX^{\top}}(t,y)\in\mathbb{Z}[t,y]$ is irreducible. By Remark 8, we may run the same argument to show that also the polynomial $\chi_{XX^{\top}}(t^{2},y)\in\mathbb{Z}[t,y]$ is irreducible. By Hilbert’s irreducibility theorem, we find infinitely many specifications of $y$ such that $\chi_{XX^{\top}}(t^{2})\in\mathbb{Z}[t]$ is irreducible of degree $2d$. The leading eigenvalue $\mu$ of $XX^{\top}$ is a root of a characteristic polynomial $\chi_{XX^{\top}}(t)$, so $\chi_{XX^{\top}}(t^{2})\in\mathbb{Z}[t]$ is the minimal polynomial of $\sqrt{\mu}$. Hence, the multicurve bipartite degree of $\alpha$ and $\beta$ equals $2d$.